Find equation of the lines containing the given pair of points..
m=2/3(5,-7)

Answer by mananth(16946)

m=2/3(5,-7)
..
the general equation is (y-y1)=m(x-x1)
y-(-7)=2/3(x-5)
y+7 = 2/3x -10/3
y = 2/3 *x -10/3 -7
y= 2/3*x -31/3
...
m.ananth@hotmail.ca